OpenStudy (anonymous):

medal=fan What is 7,200,000 written in scientific notation? A. B. C. D.

OpenStudy (anonymous):

um

OpenStudy (anonymous):

@felavin

OpenStudy (anonymous):

72x10^5

OpenStudy (anonymous):

\[7.2*10^6\]

OpenStudy (anonymous):

ok thanks! :)

OpenStudy (anonymous):

why

OpenStudy (anonymous):

@tom982

OpenStudy (anonymous):

Because 1,000,000=10^6, and 7,200,000=7.2*1,000,000=7.2*10^6

OpenStudy (anonymous):

can u help with 4 more

OpenStudy (anonymous):

Sure, but try them yourself first!

OpenStudy (jhannybean):

 that's how I imagine it.